A natural gas tank is constructed so that the pressure remains constant. On a hot day when the temperature was 33°C, the volume of gas in the tank was determined to be 3000.0 L. What would the volume be on a warm day when the temperature is 11°C?


1
Expert's answer
2021-03-09T06:07:39-0500

V1 = 3000.0 L

T1 = 33°C + 273 = 306 K

P = const

V2 = unknown

T2 = 11°C + 273 = 284 K


Solution:

Since the pressure doesn't change, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the final volume, solve the equation for V2:

V2 = V1T/ T1

V2 = (3000.0 L × 284 K) / (306 K) = 2784.3137 L = 2784.3 L

V2 = 2784.3 L


Answer: V2 = 2784.3 L.